Grade 12Mechanics

Two blocks of masses 2 kg and 4 kg are hanging with the help of massless string passing over an ideal Pulley inside in elevator. The elevator is moving upward with an acceleration g/2. The tension in the string connected between the blocks will beA..... 40 NB..... 60 NC..... 80 ND..... 20 N

Anshuman Mishra
9 Years agoGrade 12

1 Answer

Diya Chakraborty
8 Years ago
4(g+g/2)-t=4a...(1)t-2(g+g/2)=2a...(2)Therefore,t-3g=2a=>t=3g+2aSolving eqn 1from eqn (2),4(g+g/2)-(3g+2a)=4a=>3g=6a=>a=g/2Hence,t=3g+2(g/2)t=40N